# 70% Size, 100% Accuracy: Lossless LLM Compression for Efficient GPU Inference via Dynamic-Length Float

### v2

@@ Appendix F DFloat11 Compression Time
Table 5: System specifications of servers used for experiments.
GPU GPU Memory CPU CPU Memory
Server 1 NVIDIA RTX A5000 24564MiB AMD EPYC 7513 32-Core 504GB
Server 2 NVIDIA A100 40960MiB AMD EPYC 7742 64-Core 1.48TB
Server 3 NVIDIA Quadro RTX 8000 49152MiB AMD EPYC 7742 64-Core 1.48TB
@@
Figure 10: Comparison of average latency and throughput for token decoding between the original (BF16) models and their losslessly compressed (DF11) counterparts. The BF16 and DF11 models are run on the same GPU configurations, with Flash Attention [6] turned on for both methods.
Appendix H Impact of Lossy Quantization
@@ I.1 The Dual Lookup Table Approach
Huffman decoding can be performed by traversing the Huffman tree: starting from the root, each bit of the encoded bitstream determines the branch to follow, and the symbol is fully decoded upon reaching a leaf node. While this bit-by-bit traversal is conceptually simple, it is inefficient in practice. Each branching decision depends on the previous one, leading to frequent memory accesses and conditional jumps. This pattern is especially problematic on GPUs, where it causes branch divergence and limits instruction-level parallelism. A widely adopted alternative is lookup-table-based decoding [44], which flattens the Huffman tree into two compact lookup tables. This enables decoding of each symbol using just two array lookups and a bit shift, significantly improving throughput.

### v3

@@ Appendix D Pseudo-code of the GPU kernel for DFloat11 Decompression
Table 4: System specifications of servers used for experiments.
GPU GPU Memory CPU CPU Memory
Server 1 NVIDIA RTX A5000 24564MiB AMD EPYC 7513 32-Core 504GB
Server 2 NVIDIA A100 40960MiB AMD EPYC 7742 64-Core 1.48TB
Server 3 NVIDIA Quadro RTX 8000 49152MiB AMD EPYC 7742 64-Core 1.48TB
@@ Appendix F DFloat11 Compression Time
Figure 10: Comparison of average latency and throughput for token decoding between the original (BF16) models and their losslessly compressed (DF11) counterparts. The BF16 and DF11 models are run on the same GPU configurations, with Flash Attention [dao2022flashattention] turned on for both methods.
@@
Appendix H Impact of Lossy Quantization
An accuracy comparison of the original and INT8-quantized Llama model is presented in table 6.
@@ I.1 The Dual Lookup Table Approach
Huffman decoding can be performed by traversing the Huffman tree: starting from the root, each bit of the encoded bitstream determines the branch to follow, and the symbol is fully decoded upon reaching a leaf node. While this bit-by-bit traversal is conceptually simple, it is inefficient in practice. Each branching decision depends on the previous one, leading to frequent memory accesses and conditional jumps. This pattern is especially problematic on GPUs, where it causes branch divergence and limits instruction-level parallelism. A widely adopted alternative is lookup-table-based decoding [huffman_gap_arrays], which flattens the Huffman tree into two compact lookup tables. This enables decoding of each symbol using just two array lookups and a bit shift, significantly improving throughput.